OpenStudy (anonymous):

will medal factor the trinomal w^2-8w-20 A:(w-10)(w+2) B:(w-4)(w+5) C:(w+4)(w-5) D:(w+10)(w-2) E: none

OpenStudy (chosenmatt):

what do you think?

OpenStudy (anonymous):

im honestly not sure with this one

OpenStudy (uri):

Factorize it. @Alice2112 Middle term breaking Find 2 number when added/subtracted is 8 and when multiplied is 20? 10*2 =20 -10+2 =-8 So, w^2-10w-2w-20,Okay?

OpenStudy (anonymous):

okay

OpenStudy (mrnood):

you have just done a very similar question look at th eanswers the two constant terms at theend must a) MULTIPLY to give -20 b) ADD to give -8 only one answer is true for both these....

OpenStudy (mrnood):

@uri people have given several answers to identical questions it's time she tried one herself

OpenStudy (anonymous):

A? orD?

OpenStudy (uri):

@MrNood Not giving her the answer. :)

OpenStudy (anonymous):

its A

OpenStudy (uri):

Right. :)